# Chaotic Noncoherent SWIPT in Multi-Functional RIS-Aided Systems

### v1

@@ I Introduction
[…]
[lines=2]With sustainability and scalability coming out as two significant challenges for implementing the fifth generation (5G) and beyond wireless technology, worldwide research initiatives are increasingly exploring innovative solutions such as simultaneous wireless information and power transfer (SWIPT) [1]. The authors in [2] experimentally demonstrate that signals with high peak-to-average-power-ratio, like chaotic signals, provide higher wireless power transfer (WPT) efficiency. The work in [3] proposes an analytical framework for chaotic WPT, which supports the above observation. The authors in [4] investigate a noncoherent differential chaos shift keying (DCSK)-based framework and propose WPT optimal transmit waveform designs. The study in [5] proposes various link selection schemes in a relay-based DCSK-SWIPT network and the authors in [6] investigate a DCSK-based multi-antenna receiver architecture for SWIPT.
Moreover, to accommodate the rising demand of future networks with flexible and reconfigurable characteristics, a new technology termed as self-sustainable reconfigurable intelligent surface (RIS) has been recently proposed [7]. A self-sustainable RIS controls the propagation environment via software-controlled metasurfaces and its energy requirement is met from the incident signal via the process of energy harvesting (EH). The authors in [8] investigate the problem of transmit waveform design in a self-sustainable RIS-assisted DCSK-based set-up. This work focuses on the reflection-only RISs, i.e., this RIS is capable of only reflecting the incident signal. Both the transmitter (Tx) and the receiver (Rx) need to be located on the same side of the RIS. In this context, the work in [9] proposes the novel concept of a simultaneously transmitting and reflecting (STAR) RIS, capable of $360^{\circ}$ radio environment implementation. But still, the RIS transmitted/reflected signal is attenuated twice, which severely attenuates signal reaching the Rx. As a solution, the authors in [10] propose a new multi-functional (MF) RIS architecture. An MF-RIS essentially harvests energy from the incident signal, which in turn, is used for simultaneous transmission and reflection of the amplified incident signals.
@@ II System Model
We consider a set-up with a single antenna Tx and two Rxs, each equipped with a single antenna. With no direct link between the Tx and both the Rxs, an MF-RIS, consisting of $N$ elements, is employed to realize the entire communication process. Since the MF-RIS is able to support complete $360^{\circ}$ coverage, we assume that the two Rxs, namely, $U_{t}$ and $U_{r}$, to be located in each of its transmission and reflection space, respectively. Also, we consider the adjacent elements of the MF-RIS to have at least half-wavelength spacing in between, i.e., the associated wireless channels are independent.
@@ II-A SR-DCSK Signals
In conventional DCSK, each information bit is characterized by two sets of equal length chaotic samples; the first set representing the reference and the other conveying information. On the contrary, the short reference DCSK (SR-DCSK) symbol has a flexible and shorter reference length. Specifically, the SR-DCSK-based Tx output during the $p$-th transmission interval is [4]
@@
=$=\!\!\begin{cases}\!\!x_{p,q},&\!\!\!q\!=\!(p-1)(\beta+\phi)+1,\dots,(p-1)\beta+p\phi,\\ \!\!d_{p}x_{p,q-\phi},&\!\!\!q\!=\!(p-1)\beta+p\phi+1,\dots,p(\beta+\phi).\end{cases}$ (1)
@@ II-C Channel Model
where $P_{\rm t}$ is the transmission power, $h_{l,n}$ and $s_{p,q-\tau_{l}}$ denote the complex channel coefficient and the delayed transmitted signal corresponding to the $l$-th path, respectively, and $w_{n}$ is the additive white Gaussian noise (AWGN) with zero mean and variance $\frac{N_{0}}{2}$. We assume that $|h_{l,n}|=\alpha_{l,n}$ is a Rayleigh random variable [5] with $\mathbb{E}\{\alpha_{l,n}^{2}\}=\Omega_{\alpha,l,n}$, $\sum\limits_{l=1}^{L_{sr}}\Omega_{\alpha,l,n}=1$ $\forall$ $n$, and identical channel statistics across all the elements, i.e., $\Omega_{\alpha,l,1}=\Omega_{\alpha,l,2}=\dots=\Omega_{\alpha,l,N}=\Omega_{\alpha,l}$ $l=1,\dots,L_{sr}$.
Similarly, with $X\in\{t,r\}$, the received power at $U_{X}$ is again attenuated by $C_{0}d_{rd_{X}}^{-\alpha_{rd_{X}}}$, with $d_{rd_{X}}$ being the distance between the MF-RIS and $U_{X}$ and $\alpha_{rd_{X}}$ is the corresponding path loss exponent. Lastly, here also, we model the channel between the MF-RIS and $U_{X}$ as a $L_{rd_{X}}$ tap channel with complex coefficient $g_{X,n,l}$ $(n=1,\dots,N_{X}\>\>l=1,\dots,L_{rd_{X}})$, where $|g_{X,n,l}|=\beta_{X,n,l}$ is Rayleigh distributed with $\mathbb{E}\{\beta_{X,n,l}^{2}\}=\Omega_{\beta,X,n,l}$, $\sum\limits_{l=1}^{L_{rd_{X}}}\Omega_{\beta,X,n,l}=1$ $\forall$ $n$, and identical channel statistics across all MF-RIS elements.
@@ III MF-RIS-aided Chaotic System Design
In this section, the SR-DCSK frame structure is considered in the context of our MF-RIS-aided network. Note that, power transfer occurs here at the EH section, while information transfer takes place through the IT and IR section, respectively.
@@ III-A Proposed System Design
Practical scenarios suggest that the largest path delay is much smaller as compared to the reference length $\phi$ [6]. Hence, by using (3), the signal incident to the $n$-th MF-RIS patch $(n=1,\dots,N)$ can be expressed as
@@
where $\boldsymbol{y_{p,q}}\!=\!\left[y_{p,q,1},y_{p,q,2},\cdots,y_{p,q,N}\right]^{T},$ $\boldsymbol{\rm h}=\left[\sum\limits_{l=1}^{L_{sr}}h_{l,1},\right.$ $\left.\sum\limits_{l=1}^{L_{sr}}h_{l,2},\cdots,\sum\limits_{l=1}^{L_{sr}}h_{l,N}\right]^{T},$ and $\boldsymbol{\rm w}=\left[w_{1},w_{2},\cdots,w_{N}\right]^{T}.$ Since a fraction of the received signal is used for EH, IT, and IR, respectively, we rewrite $\boldsymbol{y_{p,q}}$ as
$\boldsymbol{y_{p,q}}=\left[\left(\boldsymbol{y_{p,q}^{\rm EH}}\right)^{T}\left(\boldsymbol{y_{p,q}^{\rm IT}}\right)^{T}\left(\boldsymbol{y_{p,q}^{\rm IR}}\right)^{T}\right]^{T},$ (6)
where $,\boldsymbol{y_{p,q}^{\rm EH}}\!=\!\Big{[}y_{p,q,1},\cdots,y_{p,q,N_{h}}\Big{]}^{T},\boldsymbol{y_{p,q}^{\rm IT}}\!=\!\Big{[}y_{p,q,N_{h}+1},$$\cdots,y_{p,q,N_{h}+N_{t}}\Big{]}^{T},$ and $\boldsymbol{y_{p,q}^{\rm IR}}\!=\!\Big{[}y_{p,q,N_{h}+N_{t}+1},\cdots,y_{p,q,N}\Big{]}^{T}$ denote the received signal vector at the EH, IT, and IR section, respectively.
As $\boldsymbol{y_{p,q}^{\rm IT}}$ is transmitted to $U_{t}$ by using the phase shift matrix $\Phi_{t}$, the received signal at $U_{t}$ is
$y_{p,q}^{t}$ $=\sqrt{C_{0}d_{rd_{t}}^{-\alpha_{rd_{t}}}}\boldsymbol{G_{t}}^{T}\Phi_{t}\boldsymbol{y_{p,q}^{\rm IT}}+w_{t},$ (7)
where $\boldsymbol{G_{t}}=\left[\sum\limits_{k=1}^{L_{rd_{t}}}g_{t,1,k}\sum\limits_{k=1}^{L_{rd_{t}}}g_{t,2,k}\cdots,\sum\limits_{k=1}^{L_{rd_{t}}}g_{t,N_{t},k}\right]^{T}$ and $w_{t}$ is the AWGN at $U_{t}$ with zero mean and variance $\frac{N_{0}}{2}$. By using (5) and defining $\delta_{t}=\sqrt{P_{\rm t}C_{0}^{2}d_{sr}^{-\alpha_{sr}}d_{rd_{t}}^{-\alpha_{rd_{t}}}\Upsilon_{t}},\Psi_{t}=\sqrt{C_{0}d_{rd_{t}}^{-\alpha_{rd_{t}}}\Upsilon_{t}}$, we rewrite the above expression to obtain
@@
where we have $\theta_{r,e,n}=\theta_{r,n}+\angle h_{l,N_{h}+N_{t}+n}+\angle g_{r,n,k}$ and $\theta_{r,e,n}^{w}=\theta_{r,n}+\angle g_{r,n,k}$ $\forall$ $k,l,n$, and $w_{r}$ is the AWGN at $U_{r}$ with zero mean and variance $\frac{N_{0}}{2}$.
@@
×$\times\left.\Bigg{(}\delta_{X}\sum\limits_{l=1}^{L_{sr}}\sum\limits_{k=1}^{L_{rd_{X}}}\sum\limits_{n=1}^{N_{X}}e^{j\theta_{e,X,n}}\alpha_{l,N_{h}+n}\beta_{X,n,k}x_{p,z}+\Psi_{X}\sum\limits_{k=1}^{L_{rd_{X}}}\sum\limits_{n=1}^{N_{X}}e^{j\theta^{w}_{e,X,n}}\beta_{X,n,k}w_{N_{h}+n}+w_{X,z}\Bigg{)}^{*}\right)\!\!.$ (11)
@@ III-B Information and Power Transfer
Both $U_{t}$ and $U_{r}$ recover the chaotic component from the transmitted frame to perform $\zeta$ partial correlations over each block of $\phi$ samples. By considering the aspect of low cross-correlation of two chaotic sequences, the decision metric $\lambda_{p}^{(X)}$ $X\in\{t,r\}$ corresponding to the $p$-th transmission interval, is obtained from (III-A) and (III-A), respectively, as expressed in (III-A). Thereafter, $\lambda_{p}^{(X)}$ is compared with a threshold and the actual transmitted data is recovered from the received signal.
At the EH section of the MF-RIS, each element is connected to an individual EH unit consisting of a diode followed by a low pass filter [11]. To enhance the EH performance, we employ a $\phi+\beta$ bit analog correlator [4] prior to each EH unit. An analog correlator results in effective signal integration over a certain time interval. The resultant output signal, corresponding to the $p$-th transmission interval, at the $n$-th element of the EH section is given by
@@
where $(a)$ follows from (4), by assuming that the noise is too small to be harvested. Therefore, based on the nonlinearities of the EH unit, the resulting output power is $P_{\rm EH}=\sum\limits_{n=1}^{N_{h}}\frac{v_{n,out}^{2}}{R_{L}}$, where $R_{L}$ is the load resistance and
$v_{n,out}=\eta_{1}\mathbb{E}\{|y_{p,n}^{\rm C}|^{2}\}+\eta_{2}\mathbb{E}\{|y_{p,n}^{\rm C}|^{4}\},$ (13)
with $\eta_{1},\eta_{2}$ being system parameters [11].
@@ IV-A BER Performance
${\rm BER}_{X}$ $=\frac{1}{2}\int\limits_{0}^{\infty}\int\limits_{0}^{\infty}{\rm erfc}\left(\left[\frac{(\beta+\phi)^{2}}{\gamma_{X,0}\Lambda_{X,1}\beta}\left(\frac{1}{\phi}+\frac{1}{2\gamma_{X,0}\Lambda_{X,1}}\right.\right.\right.$
$\left.\left.\left.+\frac{\Psi_{X}^{2}\Lambda_{X,2}}{2\gamma_{X,0}\Lambda_{X,1}}\left(\frac{\beta+\phi}{\phi}+\frac{4\beta\gamma_{X,0}\Lambda_{X,2}}{\phi\left(\beta+\phi\right)}\right)\right)\right]^{-\frac{1}{2}}\right)$
×$\times f(\Lambda_{X,1})f(\Lambda_{X,2})d\Lambda_{X,1}d\Lambda_{X,2},$ (14)
where $X\in\{t,r\},\gamma_{X,0}=\frac{E_{\rm b}C_{0}^{2}d_{sr}^{-\alpha_{sr}}d_{rd_{X}}^{-\alpha_{rd_{X}}}\Upsilon_{X}}{N_{0}}$,
$\Lambda_{X,1}$ $=\sum\limits_{l=1}^{L_{sr}}\!\sum\limits_{k=1}^{L_{rd_{X}}}\!\!\Big{|}\!\sum\limits_{n=1}^{N_{X}}\!e^{j\theta_{X,e,n}}\alpha_{l,N_{h}+n}\beta_{X,n,k}\Big{|}^{2},$
$\Lambda_{X,2}$ $=\sum\limits_{k=1}^{L_{rd_{X}}}\!\!\Big{|}\!\sum\limits_{n=1}^{N_{X}}\!e^{j\theta_{X,e,n}^{w}}\beta_{X,n,k}\Big{|}^{2}.$
Here, ${\rm erfc}(\cdot)$ is the complementary error function and $f(\Lambda_{X,1})$ and $f(\Lambda_{X,2})$ is the probability density function of $\Lambda_{X,1}$ and $\Lambda_{X,2}$, respectively.
@@
The probability density function of the form $\Big{|}\!\sum\limits_{n=1}^{N_{X}}\!e^{j\theta_{X,e,n}}\alpha_{l,N_{h}+n}\beta_{X,n,k}\Big{|}$ is investigated in [12] for a Nakagami-$m$ fading scenario. Hence, we can evaluate $f(\Lambda_{X,1})$ and $f(\Lambda_{X,2})$ by the standard technique of transformation of random variables. We observe that, ${\rm BER}_{X}$ is a joint function of the channel characteristics, $\Phi_{X}$ (as defined in (2)), $\beta$, and $\phi$. Note that, $\Upsilon_{X}=1$ results in negligible impact of the noise at the MF-RIS on ${\rm BER}_{X}$. Taking this into account, and by using $\Upsilon_{t}=0,\Upsilon_{r}=1$, the self-sustainable RIS discussed in [8] is demonstrated to be a special case of the MF-RIS. This can also be observed from the fact that by replacing $\Upsilon_{t}=0,\Upsilon_{r}=1$ in (1) and ignoring the MF-RIS generated noise at $U_{X}$, we obtain the BER as stated in [8, Theorem 1]. Moreover, $\Upsilon_{X}$ also depends on the distance between the MF-RIS and $U_{X}$. Therefore, in general, we have $\Upsilon_{t}\neq\Upsilon_{r}$ and hence, we individually obtain the BER for $U_{t}$ and $U_{r}$.
IV-B Average EH Performance
By using the following proposition, we look into the performance of the EH section of the MF-RIS, which consists of $N_{h}$ elements.
Proposition 2.
If the EH section of the MF-RIS consists of $N_{h}$ elements, the total harvested power is evaluated as
@@ IV-B Average EH Performance
Proof.
The Rayleigh fading channel is a special case of the generalized Nakagami-$m$ fading channel, with $m=1$. Therefore, by using this along with the fact that $\Gamma(1.5)=\frac{\sqrt{\pi}}{2}$ in [8, Theorem 3], we conclude the proof. ∎
IV-C Transmit-Reflect-Harvest Trade-off Characterization
For a SR-DCSK frame of length $\beta+\phi$, the total energy consumption of the MF-RIS is
@@ IV-C Transmit-Reflect-Harvest Trade-off Characterization
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!+\xi\left(P_{\rm t}C_{0}d_{sr}^{-\alpha_{sr}}\left(\Upsilon_{t}\!\sum\limits_{l=1}^{L_{sr}}\!\sum\limits_{n=1}^{N_{t}}\!\alpha_{l,N_{h}+n}^{2}\right.\right.$
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\left.\left.+\Upsilon_{r}\!\sum\limits_{l=1}^{L_{sr}}\!\sum\limits_{n=1}^{N_{r}}\!\alpha_{l,N_{h}+N_{t}+n}^{2}\right)\!+\!L_{sr}\frac{N_{0}}{2}\left(\Upsilon_{t}N_{t}+\Upsilon_{r}N_{r}\right)\right)\!.$
@@
From (1), we can state that for a given set of system parameters, ${\rm BER}_{X}$ is a function of $N_{X}$. Accordingly, from Proposition 2, we obtain a lower bound on $N_{h}$ as
@@
We observe that, with the other parameters remaining constant, $E_{\rm req}$ increases with $\Upsilon_{t}$ and $\Upsilon_{r}$. Therefore, even with identical choice of the $(N_{t},N_{r})$ combination, different $\Upsilon_{t}$ and $\Upsilon_{r}$ yield different $N_{h}^{\min}$. Moreover, the above bound inherently assumes $N_{h}^{\min}\leq N-(N_{t}+N_{r})$, which may not always hold. In such a scenario, an appropriate choice of transmit waveform parameters for a fixed $(N_{t},N_{r},N_{h})$ combination becomes extremely crucial. Accordingly, for a fixed $(N_{t},N_{r},N_{h})$ combination, we investigate the role of transmit waveform parameters on the BER-harvested power trade-off. Note that, to the best of our knowledge, no such investigation exists in the literature for a chaotic MF-RIS-aided system. Hence, based on the definition of the ‘success rate’ ${\rm SR}_{X}=1-{\rm BER}_{X}$ [6], we propose the generalized ${\rm SR}-P_{\rm EH}$ region definition as

### v2

@@ I Introduction
With sustainability and scalability coming out as two significant challenges for implementing the fifth generation (5G) and beyond wireless technology, worldwide research initiatives are increasingly exploring innovative solutions such as simultaneous wireless information and power transfer (SWIPT) [1]. The authors in [2] experimentally demonstrate that signals with high peak-to-average-power-ratio, such as chaotic signals, provide higher wireless power transfer (WPT) efficiency. Moreover, the work in [3] proposes an analytical framework for chaotic WPT, which supports the above observation. Furthermore, the authors in [4] investigate a noncoherent differential chaos shift keying (DCSK)-based framework and propose WPT optimal transmit waveform designs. Hence, to exploit this additional advantage for both information and power transfer, we choose DCSK-based waveforms to explore its implications for SWIPT. The study in [5] proposes various link selection schemes in a relay-based DCSK-SWIPT network and the authors in [6] investigate a DCSK-based multi-antenna receiver architecture for SWIPT.
Moreover, to accommodate the rising demand of future networks with flexible and reconfigurable characteristics, a new technology termed as reconfigurable intelligent surface (RIS) has been recently proposed [7, 8]. A self-sustainable RIS [9] controls the propagation environment via software-controlled metasurfaces and its energy requirement is met from the incident signal via the process of energy harvesting (EH). The authors in [10] investigate the problem of transmit waveform design in a self-sustainable RIS-assisted DCSK-based set-up. This work focuses on the reflection-only RISs, i.e., this RIS is capable of only reflecting the incident signal. Both the transmitter (Tx) and the receiver (Rx) need to be located on the same side of the RIS. In this context, the work in [11] proposes the novel concept of a simultaneously transmitting and reflecting (STAR) RIS, capable of $360^{\circ}$ radio environment implementation. But still, the RIS transmitted/reflected signal is attenuated twice, which severely attenuates signal reaching the Rx. As a solution, the work in [12] proposes a new multi-functional (MF) RIS, which comprises of an array of active elements. Each array element exhibits a three-layer structure, which enables multidimensional and full-space signal manipulation. An MF-RIS essentially harvests energy from the incident signal, which in turn, is used for simultaneous transmission and reflection of the amplified incident signals.
@@ II System Model
We consider a set-up with a single antenna Tx and two Rxs, each equipped with a single antenna. With no direct link between the Tx and both the Rxs, an MF-RIS, consisting of $N$ elements, is employed to realize the entire communication process. Since the MF-RIS is able to support complete $360^{\circ}$ coverage, we assume that the two Rxs, namely, $U_{t}$ and $U_{r}$, to be located in each of its transmission and reflection space, respectively. Also, we consider the adjacent elements of the MF-RIS to have at least half-wavelength spacing, i.e., the associated wireless channels are independent.
@@ II-A SR-DCSK Signals
In conventional DCSK, each information bit is characterized by two sets of equal length chaotic samples; the first set represents the reference and the other conveys information. On the contrary, the short reference DCSK (SR-DCSK) symbol has a flexible and shorter reference length, i.e., the SR-DCSK-based Tx output during the $p$-th transmission interval is [4]
@@
$=\!\!\begin{cases}\!\!x_{p,q},&\!\!\!q\!=\!(p-1)(\beta+\phi)+1,\dots,(p-1)\beta+p\phi,\\ \!\!d_{p}x_{p,q-\phi},&\!\!\!q\!=\!(p-1)\beta+p\phi+1,\dots,p(\beta+\phi),\end{cases}$ (1)
@@ II-C Channel Model
where $P_{\rm t}$ is the transmission power, $h_{l,n}$ and $s_{p,q-\tau_{l}}$ denote the complex channel coefficient and the delayed transmitted signal corresponding to the $l$-th path, respectively, and $w_{n}$ is the additive white Gaussian noise (AWGN) with zero mean and variance $\frac{N_{0}}{2}$. We assume that $|h_{l,n}|=\alpha_{l,n}$ is a Rayleigh random variable [5] with $\mathbb{E}\{\alpha_{l,n}^{2}\}=\Omega_{\alpha,l,n}$, $\sum_{l=1}^{L_{sr}}\Omega_{\alpha,l,n}=1$ $\forall$ $n$, and identical channel statistics across all the elements, i.e., $\Omega_{\alpha,l,1}=\Omega_{\alpha,l,2}=\dots=\Omega_{\alpha,l,N}=\Omega_{\alpha,l}$ $l=1,\dots,L_{sr}$.
Similarly, with $X\in\{t,r\}$, the received power at $U_{X}$ is again attenuated by $C_{0}d_{rd_{X}}^{-\alpha_{rd_{X}}}$, with $d_{rd_{X}}$ being the distance between the MF-RIS and $U_{X}$ and $\alpha_{rd_{X}}$ is the corresponding path loss exponent. Lastly, here also, we model the channel between the MF-RIS and $U_{X}$ as a $L_{rd_{X}}$ tap channel with complex coefficient $g_{X,n,l}$ $(n=1,\dots,N_{X}\>\>l=1,\dots,L_{rd_{X}})$, where $|g_{X,n,l}|=\beta_{X,n,l}$ is Rayleigh distributed with $\mathbb{E}\{\beta_{X,n,l}^{2}\}=\Omega_{\beta,X,n,l}$, $\sum_{l=1}^{L_{rd_{X}}}\Omega_{\beta,X,n,l}=1$ $\forall$ $n$, and identical channel statistics across all MF-RIS elements.
@@ III MF-RIS-aided Chaotic System Design
In this section, the SR-DCSK frame structure is considered in the context of our MF-RIS-aided network [^1] This work can also be generalized in the context of the newly proposed stacked intelligent metasurfaces, which is essentially composed of stacked multi-layer reconfigurable surfaces [13].. Note that, power transfer occurs here at the EH section, while information transfer takes place through the IT and IR section, respectively.
@@ III-A Proposed System Design
Practical scenarios suggest that the largest path delay is much smaller as compared to the reference length $\phi$ [5, 6]. Hence, by using (3), the signal incident to the $n$-th MF-RIS patch $(n=1,\dots,N)$ can be expressed as
@@
where $\boldsymbol{y_{p,q}}\!=\!\left[y_{p,q,1},y_{p,q,2},\cdots,y_{p,q,N}\right]^{T},$ $\boldsymbol{\rm h}=\left[\sum_{l=1}^{L_{sr}}h_{l,1},\sum_{l=1}^{L_{sr}}h_{l,2},\cdots,\sum_{l=1}^{L_{sr}}h_{l,N}\right]^{T}\!$, and $\boldsymbol{\rm w}=\left[w_{1},w_{2},\right$. $\left.\cdots,w_{N}\right]^{T}\!$. Since a fraction of $\boldsymbol{y_{p,q}}$ is used for EH, IT, and IR, respectively, we rewrite (5) as $\boldsymbol{y_{p,q}}=\left[\left(\boldsymbol{y_{p,q}^{\rm EH}}\right)^{T}\left(\boldsymbol{y_{p,q}^{\rm IT}}\right)^{T}\left(\boldsymbol{y_{p,q}^{\rm IR}}\right)^{T}\right]^{T},$ where $\boldsymbol{y_{p,q}^{\rm EH}}=$ $\Big[y_{p,q,1},\cdots,y_{p,q,N_{h}}\Big]^{T},\boldsymbol{y_{p,q}^{\rm IT}}\!=\!\Big[y_{p,q,N_{h}+1},\cdots,$ $y_{p,q,N_{h}+N_{t}}\Big]^{T}$, and $\boldsymbol{y_{p,q}^{\rm IR}}\!=\!\Big[y_{p,q,N_{h}+N_{t}+1},\cdots,y_{p,q,N}\Big]^{T}$ denote the received signal vector at the EH, IT, and IR section, respectively.
As $\boldsymbol{y_{p,q}^{\rm IT}}$ is transmitted to $U_{t}$ by using the phase shift matrix $\Phi_{t}$, the received signal at $U_{t}$ is $y_{p,q}^{t}=\sqrt{C_{0}d_{rd_{t}}^{-\alpha_{rd_{t}}}}\boldsymbol{g_{t}}^{T}\Phi_{t}\boldsymbol{y_{p,q}^{\rm IT}}+w_{t},$ where $\boldsymbol{g_{t}}=$ $\left[\sum_{k=1}^{L_{rd_{t}}}g_{t,1,k}\sum_{k=1}^{L_{rd_{t}}}g_{t,2,k}\cdots,\sum_{k=1}^{L_{rd_{t}}}g_{t,N_{t},k}\right]^{T}$ and $w_{t}$ is the AWGN at $U_{t}$ with zero mean and variance $\frac{N_{t,0}}{2}$. By using (5) and defining $\delta_{t}=\sqrt{P_{\rm t}C_{0}^{2}d_{sr}^{-\alpha_{sr}}d_{rd_{t}}^{-\alpha_{rd_{t}}}\Upsilon_{t}}$ and $\Psi_{t}=$ $\sqrt{C_{0}d_{rd_{t}}^{-\alpha_{rd_{t}}}\Upsilon_{t}}$, we rewrite the above expression to obtain
@@
where we have $\theta_{r,e,n}=\theta_{r,n}+\angle h_{l,N_{h}+N_{t}+n}+\angle g_{r,n,k}$ and $\theta_{r,e,n}^{w}=\theta_{r,n}+\angle g_{r,n,k}$ $\forall$ $k,l,n$, and $w_{r}$ is the AWGN at $U_{r}$ with zero mean and variance $\frac{N_{r,0}}{2}$.
@@
$\times\left.\Bigg(\delta_{X}\sum\limits_{l=1}^{L_{sr}}\sum\limits_{k=1}^{L_{rd_{X}}}\sum\limits_{n=1}^{N_{X}}e^{j\theta_{e,X,n}}\alpha_{l,N_{h}+n}\beta_{X,n,k}x_{p,z}+\Psi_{X}\sum\limits_{k=1}^{L_{rd_{X}}}\sum\limits_{n=1}^{N_{X}}e^{j\theta^{w}_{e,X,n}}\beta_{X,n,k}w_{N_{h}+n}+w_{X,z}\Bigg)^{*}\right)\!\!.$ (9)
@@ III-B Information and Power Transfer
Both $U_{t}$ and $U_{r}$ recover the chaotic component from the transmitted frame to perform $\zeta$ partial correlations over each block of $\phi$ samples. By considering the aspect of low cross-correlation of two chaotic sequences [10], the decision metric $\lambda_{p}^{(X)}$ $X\in\{t,r\}$ corresponding to the $p$-th transmission interval, is obtained from (III-A) and (III-A), respectively, as expressed in (III-A), where denotes the complex conjugate. Thereafter, $\lambda_{p}^{(X)}$ is compared with a threshold and the actual transmitted data is recovered from the received signal.
At the EH section of the MF-RIS, each element is connected to an individual EH unit consisting of a diode followed by a low pass filter [14]. To enhance the EH performance, we employ a $\phi+\beta$ bit analog correlator [4] prior to each EH unit. An analog correlator is a low complexity yet very effective signal integrator, which simply sums the received signal over a certain time interval. The resultant output signal, corresponding to the $p$-th transmission interval, at the $n$-th element of the EH section is given by
@@
where $(a)$ follows from (4), by assuming that the noise is too small to be harvested. Therefore, based on the nonlinearities of the EH unit, the resulting output power is $P_{\rm EH}=\sum_{n=1}^{N_{h}}\frac{v_{n,out}^{2}}{R_{L}}$, where $R_{L}$ is the load resistance and $v_{n,out}=\eta_{1}\mathbb{E}\{|y_{p,n}^{\rm C}|^{2}\}+\eta_{2}\mathbb{E}\{|y_{p,n}^{\rm C}|^{4}\},$ with $\eta_{1},\eta_{2}$ being system parameters [14].
@@ IV Waveform Design Investigating Transmit-Reflect-Harvest Trade-off
[…]
$\!\!\!\!\!\!\!\!\!\!\!\!\left.\left.\left.+\frac{\Psi_{X}^{2}\Lambda_{X,2}}{2\gamma_{X,0}\Lambda_{X,1}}\left(\frac{N_{0}}{N_{X,0}}\right)\left(\frac{\beta+\phi}{\phi}+\frac{4\beta\gamma_{X,0}\Lambda_{X,1}}{\phi\left(\beta+\phi\right)}\right)\right)\right]^{-\frac{1}{2}}\right)$
$\!\!\!\!\!\!\!\!\!\!\times f(\Lambda_{X,1})f(\Lambda_{X,2})d\Lambda_{X,1}d\Lambda_{X,2},$ (11)
where $X\in\{t,r\},\gamma_{X,0}=\frac{E_{\rm b}C_{0}^{2}d_{sr}^{-\alpha_{sr}}d_{rd_{X}}^{-\alpha_{rd_{X}}}\Upsilon_{X}}{N_{t,0}}$, $\Lambda_{X,1}=\sum_{l=1}^{L_{sr}}\!\sum_{k=1}^{L_{rd_{X}}}\!\!\Big|\!\sum_{n=1}^{N_{X}}\!e^{j\theta_{X,e,n}}\alpha_{l,N_{h}+n}\beta_{X,n,k}\Big|^{2}$, and $\Lambda_{X,2}=\sum_{k=1}^{L_{rd_{X}}}\!\!\Big|\!\sum_{n=1}^{N_{X}}\!e^{j\theta_{X,e,n}^{w}}\beta_{X,n,k}\Big|^{2}.$ Here, ${\rm erfc}(\cdot)$ is the complementary error function and $f(\Lambda_{X,1})$ and $f(\Lambda_{X,2})$ is the probability density function of $\Lambda_{X,1}$ and $\Lambda_{X,2}$, respectively.
@@
The probability density function of the form $\Big|\!\sum_{n=1}^{N_{X}}\!e^{j\theta_{X,e,n}}\alpha_{l,N_{h}+n}\beta_{X,n,k}\Big|$ is investigated in [15] for a Nakagami-$m$ fading scenario. Hence, we can evaluate $f(\Lambda_{X,1})$ and $f(\Lambda_{X,2})$ by using $m=1$ in [15, Proposition 1] followed by the standard technique of transformation of random variables. We observe that, ${\rm BER}_{X}$ is a joint function of the channel characteristics, $\Phi_{X}$ (as defined in (2)), $\beta$, and $\phi$. Note that, $\Upsilon_{X}=1$ results in negligible impact of the noise at the MF-RIS on ${\rm BER}_{X}$. Taking this into account, and by using $\Upsilon_{t}=0,\Upsilon_{r}=1$, the self-sustainable RIS discussed in [10] is demonstrated to be a special case of the MF-RIS. This can also be observed from the fact that by replacing $\Upsilon_{t}=0,\Upsilon_{r}=1$ in (1) and ignoring the MF-RIS generated noise at $U_{X}$, we obtain the BER as stated in [10, Theorem 1]. Moreover, $\Upsilon_{X}$ also depends on the distance between the MF-RIS and $U_{X}$. Therefore, in general, we have $\Upsilon_{t}\neq\Upsilon_{r}$ and hence, we individually obtain the BER for $U_{t}$ and $U_{r}$. Furthermore, (1) also demonstrates the impact of the noise generated at the MF-RIS on the BER performance. Specifically, based on the nature of ${\rm erfc}(x)$ for $x\geq 0$, we conclude that, for a given set of system parameters, an increasing $N_{0}$ translates to a deteriorating BER performance.
Next, we look into the performance of the EH section of the MF-RIS, which consists of $N_{h}$ elements. Accordingly, by using the Nakagami-$m$ parameter as unity and $\Gamma(1.5)=\frac{\sqrt{\pi}}{2}$ in [10, Theorem 3], the total harvested power is evaluated as
@@
Note that, for a SR-DCSK frame of length $\beta+\phi$, the total energy consumption of the MF-RIS is
@@
[…]
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\left.\left.+\Upsilon_{r}\!\sum\limits_{l=1}^{L_{sr}}\!\sum\limits_{n=1}^{N_{r}}\!\alpha_{l,N_{h}+N_{t}+n}^{2}\right)\!+\!\frac{L_{sr}}{2}\left(\Upsilon_{t}N_{t}N_{t,0}+\Upsilon_{r}N_{r}N_{r,0}\right)\right)\!.$
@@
From (1), we can state that for a given set of system parameters, ${\rm BER}_{X}$ is a function of $N_{X}$. Accordingly, from (IV), we obtain a lower bound on $N_{h}$ as
@@
We observe that, with the other parameters remaining constant, $E_{\rm req}$ increases with $\Upsilon_{t}$ and $\Upsilon_{r}$. Therefore, even with identical choice of the $(N_{t},N_{r})$ combination, different $\Upsilon_{t}$ and $\Upsilon_{r}$ yield different $N_{h}^{\min}$. Moreover, the above bound inherently assumes $N_{h}^{\min}\leq N-(N_{t}+N_{r})$, which may not always hold. In such a scenario, an appropriate choice of transmit waveform parameters for a fixed $(N_{t},N_{r},N_{h})$ combination becomes extremely crucial. Therefore, for a fixed $(N_{t},N_{r},N_{h})$ combination, we investigate the role of transmit waveform parameters on the BER-harvested power trade-off. Note that, to the best of our knowledge, no such investigation exists in the literature for a chaotic MF-RIS-aided system. Hence, based on the definition of the ‘success rate’ ${\rm SR}_{X}=1-{\rm BER}_{X}$ [6], we propose the generalized ${\rm SR}-P_{\rm EH}$ region definition as